When the volume is held constant, the relationship between pressure and temperature is

a. inversely proportional

b. vertical

c. quadratic

d. directly proportional
Answer #1

Answer : d) directly proportional

from ideal gas law :

P V = n R T

here R and n are constant . (n = moles , R = gas constant)

if volume is constant .

P / T = n R / V

P / T = constant

P1 / T1 = P2 / T2

Pressure and temperature , are directly proportional